Your firm has a credit rating of A. You notice that the credit spread for five​-year...

Your

firm has a credit rating of A.

You

notice

that the credit spread for

five​-year

maturity A debt is

85

basis points

left parenthesis 0.85 % right parenthesis .(0.85%).

Your

​firm's

five​-year

has​ semi-annual coupons and a coupon rate of

6​%.

You

see

that new

five​-year

Government of Canada bonds are being issued with a

YTM

of

3​%.

What should the price of

your

outstanding

five​-year

bonds​ be? Assume a par value of

​$100.
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASEe v ( 02:03 ENG 05-03-207022 BV398 x x BQ BR BS BT BUBV B W BX BY 381 382 383 BOND BA II PLUS set C/Y = P/Y = 384 385 386 387